OpenStudy (anonymous):

f(x) = x^3 + 4x^2 - x - 4 Part A: What are the factors of f(x)? Part B: What are the zeros of f(x)? Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x).

hero (hero):

HInt: x^3 + 4x^2 = x^2(x + 4) - x - 4 = -x(x + 4)